Prove that if a complex number z is an nth root of a real number a , then so
is z bar.
1
Expert's answer
2020-02-20T08:58:47-0500

We know that for every complex number which is in the form of a+ib

now, for any positive integer n let take nth root in the form of x+iy

so,

(a+ib)n = x+iy

Now use De Moivre's Theorem to extend it

so complex number , z= r("cos\theta + sin \theta" i)

which has n distinct root can be determined by "\sqrt [n]{z}=\sqrt[n]{r}(cos \alpha+i sin\alpha)"

where , "(\alpha = \theta +360k)/n)"

we can take , k=0,1,2,3..........(n-1)

so using the above relation we can get the nth root of complex number and which is in the form of z bar as, "z= r(cos\u03b8+sin\u03b8 i)"

Take conjugate of general equation , as

"\bar{z}=a-ib" In polar coordinates form it will be "\bar {z} = r(cos \theta -i sin \theta )"

If we find roots in the conjugate case and using De Moivre,s theorem

to expand the below"\sqrt [n]{z}=\sqrt[n]{r}(cos \alpha-i sin\alpha)"

we get in this case also we are getting real part as a which is same as of z

Hence proved